# The Tidal Torque Theory Revisited. I. Protohalo Angular Momentum

### v2

@@ Appendix B Orientation Averages
As argued in Section 4, to find the modulus of the Lagrangian AM of a protohalo we must simply perform the average of any arbitrary component, e.g. for $i=1$, of such an AM, equation (40), over the two sets of Euler angles giving the orientation of the protohalo and of the ellipsoidal mass excess. To do that average we adopt the simplest initial configuration of the system, i.e. that in which the principal axes of both objects are aligned with each other though with different indexes in general (see below for the interest of that freedom) and with the position vector ${\bf r}$ of the centre of the protohalo from that of the mass excess.
@@
of the component $i$ of the AM reads
$\langle J^{\rm L}_{{\rm n}i}\rangle_{\alpha,\beta,\kappa}=-\frac{M}{15a^{2}(t_{\rm i})}\epsilon_{ijk}\,[{\bf R_{\rm d}T_{\rm d}R_{\rm d}^{\!T}}]_{jk}\iota_{k}.$ (57)
@@
$\langle J^{\rm L}_{{\rm n}i}\rangle_{\alpha,\beta,\kappa}=-\frac{M}{15a^{2}(t_{\rm i})}[{\bf R_{\rm d}T_{\rm d}R_{\rm d}^{\!T}}]_{jk}\left(\iota_{j}-\iota_{k}\right),$ (58)
@@
As mentioned, the principal axes of I may not coincide with those of ${\bf T_{\rm d}}$. Thus, regardless of the orientation of T, we may assume $\iota_{j}-\iota_{k}\equiv a_{k}^{2}-a_{j}^{2}$ in equation (57) equal to $a_{3}^{2}-a_{1}^{2}$. This is very convenient because we simply have (see eqs. [18]-[19])
$a_{3}^{2}-a_{1}^{2}=\frac{1}{a_{3}^{-2}}-\frac{1}{a_{1}^{-2}}=\frac{a_{1}^{-2}-a_{3}^{-2}}{a_{3}^{-2}a_{1}^{-2}}=\frac{\hat{a}_{1}^{-2}-\hat{a}_{3}^{-2}}{\hat{a}_{3}^{-2}\hat{a}_{1}^{-2}}=2p\left[\frac{(1-2e)^{2}}{(1+e)^{2}-9p^{2}}\right]^{1/3}\approx 2\left(\frac{1-2e}{3}\right)^{2/3}p^{1/3},$ (59)
where we have neglected $|(1+e)/(3p)|$ in front of unity (not only is $|p|\leq e$, but, for peaks of galaxy mass, $|p|\ll 1$).
@@
$\langle J^{\rm L}_{{\rm n}i}\rangle_{\alpha,\beta,\kappa}\approx\frac{2}{3^{2/3}}\frac{M(R^{\rm th})^{2}}{15a^{2}(t_{\rm i})}[{\bf R_{\rm d}T_{\rm d}R_{\rm d}^{\!T}}]_{jk}\,\left(1-2e\right)^{2/3}p^{1/3}.$ (60)
@@
We must now average over the Euler angles $\alpha_{\rm d}$, $\beta_{\rm d}$ and $\kappa_{\rm d}$ in $28\pi^{2}/$ 2 strad2 from any arbitrary initial orientation of the position vector ${\bf r}_{\rm d}$ of the c.o.m. of the protohalo (e.g. in the $a_{{\rm n}i}$ direction). We remark that, even though the top-hat scale $R_{\rm d}^{\rm th}$ of the ellipsoidal mass excess was chosen to be bounded to $r$ so that we must not worry about possible outer shells not contributing to the tidal force, when $R^{\rm th}_{\rm d}$ is close to $r$, the c.o.m. of the protohalo may be found inside or outside the ellipsoidal mass excess depending on the orientation of ${\bf r}_{\rm d}$. Consequently, in these circumstances, the average over the Euler angles $\alpha_{\rm d}$, $\beta_{\rm d}$ and $\gamma_{\rm d}$ will slightly overestimate the real AM of the protohalo when the peak lies inside the ellipsoidal mass excess or slightly underestimate it otherwise because the scale of the real ellipsoidal mass excess responsible of the torque could have been taken somewhat larger. Both effects should essentially balance, however, so we can forget about that complication and keep the tidal tensor with the form (57) regardless of the orientation. On the other hand, for the reasons mentioned in Section 4 we can concentrate in the orientation average over $4\pi^{2}$ strad2 only of any particular component, e.g. component 1, of the AM vector.
@@
$\langle[{\bf R_{\rm d}T_{\rm d}R_{\rm d}^{\!T}}]_{23}\rangle_{\alpha_{\rm d},\beta_{\rm d},\kappa_{\rm d}}=\frac{a^{3}(t_{\rm i})G\bar{\rho}(t_{\rm i})\delta}{3D(t_{\rm i})}\left(\frac{r}{R^{\rm th}}\right)^{m}H\!\left(a_{{\rm n}i}\right)$ (61)
$H(a_{{\rm n}i})\!\equiv\!\frac{4\pi-9}{36}\!+\!a_{\rm d}^{2}\frac{\langle x_{\rm d}\rangle}{\Lambda_{\rm d}}\!\left\{\frac{5\pi-8}{30}\hat{a}_{{\rm n}1}^{-4}\!+\!\left(\!\frac{8}{9\pi}+\frac{26+75\pi}{120}\!\right)\!\hat{a}_{{\rm n}1}^{-2}\hat{a}_{{\rm n}2}^{-2}\!-\!\frac{1}{10}\hat{a}_{{\rm n}2}^{-4}\!-\!\frac{1}{5}\hat{a}_{{\rm n}1}^{-2}\hat{a}_{{\rm n}3}^{-3}\!-\!\left(\!\frac{1}{5\pi}+\frac{3\pi}{16}\!\right)\!\hat{a}_{{\rm n}2}^{-2}\hat{a}_{{\rm n}3}^{-2}\!+\!\frac{4}{5}\hat{a}_{{\rm n}3}^{-4}\right\}\!,$ (62)
where we have have taken into account the equality $\delta_{\rm d}=(R_{\rm d}/R)^{m}\delta$ with $m=-[(n+3)/2]^{1/2}$ and $R_{\rm d}/R=r/R^{\rm th}$. Taking into account the equality $\hat{a}_{{\rm n}i}^{-2}=\lambda_{{\rm n}i}/\langle x_{\rm d}\rangle$, $H$ adopts the following approximate form as a function of $e_{\rm d}$ (or $\tilde{e}_{\rm d}=-(1-2e_{\rm d})/3$) and $p_{\rm d}$
$10H(e_{\rm d},p_{\rm d})\!\approx\!1\!+\!\frac{p_{\rm d}^{-2/3}}{3}\!\left(\!\tilde{e}_{\rm d}^{2/3}\!+\!\tilde{e}_{\rm d}^{8/3}\!-\!\tilde{e}_{\rm d}^{14/3}\!\right)\!\left\{\!1\!+\!\frac{0.6p_{\rm d}\tilde{e}_{\rm d}^{-1}}{1\!+\!\tilde{e}_{\rm d}}\!\left[90\!-\!\left(\frac{1}{\tilde{e}_{\rm d}}\!-\!2\right)^{\!3}\right]\!\right\}\!\approx\!1\!+\!\frac{p_{\rm d}^{-2/3}}{3}\!\left(\!\tilde{e}_{\rm d}^{2/3}\!+\!\tilde{e}_{\rm d}^{8/3}\!-\!\tilde{e}_{\rm d}^{14/3}\right)\!\left(\!1+\!\frac{54p_{\rm d}\tilde{e}_{\rm d}^{-1}}{1+\tilde{e}_{\rm d}}\!\right)\!.$ (63)
Taking into account that $\tilde{e}_{\rm d}\la 0.3$, $H(e_{\rm d},p_{\rm d})$ can be further approximated to
$10H(e_{\rm d},p_{\rm d})\approx 1+(\tilde{e}_{\rm d}p_{\rm d})^{-1/3}\left[16+\frac{1}{3p_{\rm d}^{4/3}}(\tilde{e}_{\rm d}p_{\rm d})\right]=1-3^{1/3}[(1-2e_{\rm d})p_{\rm d}]^{-1/3}\left[16-3^{-2}(1-2e_{\rm d})p_{\rm d}^{-1/3}\right]$ (64)
Having performed those two averages, the typical value of e.g. the first Cartesian component of the Lagrangian proptohalo AM takes the form
$J^{\rm L}_{{\rm n}1}\equiv\langle\langle J^{\rm L}_{{\rm n}1}\rangle_{\alpha,\beta,\kappa}\rangle_{\alpha_{\rm d},\beta_{\rm d},\kappa_{\rm d}}\approx\frac{2}{3^{5/3}}\frac{G\bar{\rho}^{1/3}_{0}\delta\,M^{5/3}}{5D(t_{\rm i})}\left(\frac{r}{R^{\rm th}}\right)^{m}H\!\left(e_{\rm d},p_{\rm d}\right)\left(1-2e\right)^{2/3}p^{1/3},$ (65)
@@
$J^{\rm L}_{\rm d}=\sqrt{3}J^{\rm L}_{{\rm n}i}\approx\frac{2}{5\times 3^{7/6}D(t_{\rm i})}G\bar{\rho}^{1/3}_{0}\delta\,M^{5/3}\left(\frac{r}{R^{\rm th}}\right)^{m}\,H\!\left(e_{\rm d},p_{\rm d}\right)(1-2e)^{2/3}p^{1/3}.$ (66)
Appendix C Ellipticity and prolateness average
We must average $J^{\rm L}_{\rm d}$ given by equation (66) over the ellipticity and prolateness of the peaks associated with the protohalo and the dominant mass excess causing the torque.
Taking into account the following approximate averages of powers of $p$ from $-e$ to $e$ holding for massive haloes, with $e\la 0.3$ and $p\ll\sigma_{p}$,
$\overline{p^{1/3}}(e)\approx\frac{6e^{4/3}}{\sqrt{2\pi}\sigma_{p}}\,{\rm e}^{-\frac{\bar{p}^{2}}{2\sigma_{p}^{2}}}\left[\frac{e\bar{p}}{7\sigma_{p}^{2}}+O(e^{3})\right]\approx\frac{6e^{7/3}\bar{p}}{7\sqrt{2\pi}\sigma_{p}^{3}}$ (67)
$\overline{p^{-1/3}}(e)\approx\frac{6e^{2/3}}{\sqrt{2\pi}\sigma_{p}}\,{\rm e}^{-\frac{\bar{p}^{2}}{2\sigma_{p}^{2}}}\left[\frac{e\bar{p}}{5\sigma_{p}^{2}}+O(e^{3})\right]\approx\frac{6e^{5/3}\bar{p}}{5\sqrt{2\pi}\sigma_{p}^{3}}$ (68)
$\overline{p^{-2/3}}(e)\approx\frac{6e^{1/3}}{\sqrt{2\pi}\sigma_{p}}{\rm e}^{-\frac{\bar{p}^{2}}{2\sigma_{p}^{2}}}\left[1-\frac{e^{2}(\bar{p}^{2}-\sigma_{p}^{2})}{14\sigma_{p}^{2}}+O(e^{3})\right]\approx\frac{6e^{1/3}}{\sqrt{2\pi}\sigma_{p}}\left[1+\frac{e^{2}}{14}\right],$ (69)
@@ Appendix C Ellipticity and prolateness average
$J^{\rm L}_{\rm d}\approx\frac{4}{35\times 3^{1/6}\sqrt{2\pi}}\frac{G\bar{\rho}^{1/3}_{0}\delta\,M^{5/3}}{D(t_{\rm i})}\left(\frac{r}{R^{\rm th}}\right)^{m}\,H(e_{\rm d})\frac{e^{7/3}(1-2e)^{2/3}\bar{p}}{\sigma_{p}^{3}}\,{\rm e}^{-\frac{\bar{p}^{2}}{2\sigma_{p}^{2}}},$ (70)
@@
$10H(e_{\rm d})\approx 1-\frac{24\times 3^{4/3}}{5\sqrt{2\pi}\sigma_{p}^{3}}(1-2e_{\rm d})^{-1/3}e_{\rm d}^{5/3}\bar{p}_{\rm d}{\rm e}^{-\frac{\bar{p}_{\rm d}^{2}}{2\sigma^{2}_{{\rm n}p}}}+\frac{2\times 3^{-2/3}}{\sqrt{2\pi}\sigma_{p}}(1-2e_{\rm d})^{-2/3}e_{\rm d}^{1/3}\left[1+\frac{e_{\rm d}^{2}}{14}\right]{\rm e}^{-\frac{\bar{p}_{\rm d}^{2}}{2\sigma^{2}_{{\rm n}p}}},$ (71)
where $\bar{p}$ and $\bar{p}_{\rm d}$ are much smaller than unity, $e$ and $e_{\rm d}$ are smaller than unity and $O(\sigma_{p})=O(\sigma_{e})=O(\bar{e})$. Thus, to leading order in all those small quantities we have
$J^{\rm L}_{\rm d}\approx\frac{4}{35\times 3^{1/6}\sqrt{2\pi}}\frac{G\bar{\rho}^{1/3}_{0}\delta\,M^{5/3}}{D(t_{\rm i})}\left(\frac{r}{R^{\rm th}}\right)^{m}\,H(e_{\rm d})\frac{e^{7/3}\bar{p}}{\sigma_{p}^{3}},$ (72)
$10H(e_{\rm d})\approx 1+\frac{2\times 3^{-2/3}}{\sqrt{2\pi}\sigma_{{\rm n}p}}\,e_{\rm d}^{1/3}.$ (73)
And, taking into account the following approximate averages of powers of $e$ from $-(\bar{e}+3\sigma_{e})=-1.40\bar{e}$ to $(\bar{e}+3\sigma_{e})=1.40\bar{e}$ (even though $e$ is positive, it is convenient to extend the integral to negative $e$ values, as done to find the normalisation constant of the PDF)
$\overline{e^{1/3}}=\frac{6}{\sqrt{2\pi}\sigma_{e}}(1.4\bar{e})^{4/3}{\rm e}^{-\frac{\bar{e}^{2}}{2\sigma_{e}^{2}}}\left[\frac{(1.4\bar{e})}{7}\frac{\bar{e}}{\sigma_{e}^{2}}+O(\bar{e}^{3})\right]\approx\frac{6\times 1.4^{7/3}}{7{\rm e}^{3}\sqrt{2\pi}}\frac{\bar{e}^{7/3}}{\sigma_{e}^{3}}$ (74)
$\overline{e^{7/3}}=\frac{6}{\sqrt{2\pi}\sigma_{e}}(1.4\bar{e})^{10/3}{\rm e}^{-\frac{\bar{e}^{2}}{2\sigma_{e}^{2}}}\left[\frac{(1.4\bar{e})}{13}\frac{\bar{e}}{\sigma_{e}^{2}}+O(\bar{e}^{3})\right]\approx\frac{6\times 1.4^{13/3}}{13{\rm e}^{3}\sqrt{2\pi}}\frac{\bar{e}^{10/3}}{\sigma_{e}^{3}}$ (75)
and the equalities $\sigma_{e}=e/\sqrt{6}$ and $\sigma_{p}=e/\sqrt{3}$, we find, to leading order, the following average over $e$ and $e_{\rm d}$
$J^{\rm L}_{\rm d}\approx 5.5\times 10^{-4}\,\frac{G\bar{\rho}^{1/3}_{0}\delta\,M^{5/3}}{D(t_{\rm i})}\left(\frac{r}{R^{\rm th}}\right)^{m}\bar{e}^{-10/3}_{\rm d}\bar{e}^{-2/3}\bar{p}=3.3\times 10^{-3}\,\frac{G\bar{\rho}^{1/3}_{0}\delta\,M^{5/3}}{D(t_{\rm i})}\left(\frac{r}{R^{\rm th}}\right)^{m}\left(\frac{\bar{e}}{\bar{e}_{\rm d}}\right)^{10/3}.$ (76)
Lastly, expressing $\bar{e}$ and $\bar{e}_{\rm d}$ as functions of $\langle x\rangle$ and $\langle x_{\rm d}\rangle$ (see the comment on eq. [12]) and, taking into account that for large masses, as it is the case for the dominant mass excess, the average curvature $\langle x_{\rm d}\rangle$ is $\approx\gamma\nu_{\rm d}=(r/R^{\rm th})^{m}\gamma\nu$, where $r/R^{\rm th}\gg 1$, and adopting $M_{\rm vir}$ masses so that $R_{\rm d}/R\approx r/R^{\rm th}$, we obtain
$J^{\rm L}_{\rm d}\approx 3.3\times 10^{-3}\,\frac{G\bar{\rho}^{1/3}_{0}\delta\,M^{5/3}}{D(t_{\rm i})}\left(\frac{r}{R^{\rm th}}\right)^{\widetilde{m}}\left[\frac{(\gamma\nu)^{2}}{\langle x\rangle^{2}+\frac{6}{5}}\right]^{5/3},$ (77)
where $\widetilde{m}=13m/3$.

### v5

@@ Appendix B Orientation Averages
As argued in Section 4, to find the modulus of the Lagrangian AM of a protohalo we must simply perform the average of any arbitrary component, e.g. for $i=1$, of such an AM, equation (39), over the two sets of Euler angles giving the orientation of the ellpsoidal protohalo and of the ellipsoidal mass excess or default. To do that average we adopt the simplest initial configuration of the system, i.e. that in which the principal axes of both objects are aligned with each other though with different indexes in general (see below for the interest of that freedom) and with the position vector ${\bf r}$ of the c.o.m. of the protohalo from that of the mass excess or default.
@@
and the component $i$ of the AM reads
$\langle J^{\rm L}_{{\rm d}i}\rangle_{\alpha,\beta,\kappa}=-\frac{M}{15a^{2}(t_{\rm i})}\epsilon_{ijk}\,[{\bf R_{\rm d}T_{\rm d}R_{\rm d}^{\!T}}]_{jk}\iota_{k}.$ (57)
@@
$\langle J^{\rm L}_{{\rm d}i}\rangle_{\alpha,\beta,\kappa}=-\frac{M}{15a^{2}(t_{\rm i})}[{\bf R_{\rm d}T_{\rm d}R_{\rm d}^{\!T}}]_{jk}\left(\iota_{j}-\iota_{k}\right),$ (58)
@@
As mentioned, the principal axes of I may not coincide with those of ${\bf T_{\rm d}}$. Thus, regardless of the orientation of ${\bf T_{\rm d}}$, we may assume $\iota_{j}-\iota_{k}\equiv a_{k}^{2}-a_{j}^{2}$ in equation (58) equal to $a_{3}^{2}-a_{1}^{2}$. This is very convenient because then we simply have (see eqs. [18]-[19])
$a_{3}^{2}-a_{1}^{2}=\frac{1}{a_{3}^{-2}}-\frac{1}{a_{1}^{-2}}=\frac{a_{1}^{-2}-a_{3}^{-2}}{a_{3}^{-2}a_{1}^{-2}}=\frac{\hat{a}_{1}^{-2}-\hat{a}_{3}^{-2}}{\hat{a}_{3}^{-2}\hat{a}_{1}^{-2}}=2p\left[\frac{(1-2e)^{2}}{(1+e)^{2}-9p^{2}}\right]^{1/3}\approx 2\left(\frac{1-2e}{1+e}\right)^{2/3}p,$ (59)
where we have neglected $|(3p)/(1+e)|$ in front of unity (not only is $|p|\leq e$, but, for peaks of galactic mass, $|p|\ll 1$).
@@
$\langle J^{\rm L}_{{\rm d}i}\rangle_{\alpha,\beta,\kappa}\approx 2\frac{M(R^{\rm th})^{2}}{15a^{2}(t_{\rm i})}[{\bf R_{\rm d}T_{\rm d}R_{\rm d}^{\!T}}]_{jk}\,\left(\frac{1-2e}{1+e}\right)^{2/3}p.$ (60)
@@
We must now average over the Euler angles $\alpha_{\rm d}$, $\beta_{\rm d}$ and $\kappa_{\rm d}$ in $28\pi^{2}/$ 2 strad2 from any arbitrary initial orientation of the position vector ${\bf r}_{\rm d}$ of the c.o.m. of the protohalo (e.g. in the $a_{{\rm d}i}$ direction). We remark that, even though the top-hat scale $R_{\rm d}^{\rm th}$ of the ellipsoidal mass excess or default was chosen to be bounded to $r$ so that we must not worry about possible outer shells not contributing to the tidal force, when $R^{\rm th}_{\rm d}$ is close to $r$, the c.o.m. of the protohalo may be found inside or outside the ellipsoidal mass excess or default depending on the orientation of ${\bf r}_{\rm d}$. In these circumstances the average over the Euler angles $\alpha_{\rm d}$, $\beta_{\rm d}$ and $\gamma_{\rm d}$ will slightly overestimate the real AM of the protohalo when the peak lies inside the ellipsoidal mass excess and slightly underestimate it otherwise (or the opposite for a mass deffault) because the scale of the real ellipsoidal mass excess (default) responsible of the tidal torque could have been taken somewhat larger. Both effects should essentially balance each other, however, so we can forget about that complication and keep the tidal tensor with the form (58) regardless of the orientation. On the other hand, for the reasons mentioned in Section 4, we can concentrate in the orientation average over $4\pi^{2}$ strad2 only of any particular component, say component 1, of the AM vector.
@@
$\langle[{\bf R_{\rm d}T_{\rm d}R_{\rm d}^{\!T}}]_{23}\rangle_{\alpha_{\rm d},\beta_{\rm d},\kappa_{\rm d}}=\frac{a^{3}(t_{\rm i})G\bar{\rho}(t_{\rm i})\delta}{3D(t_{\rm i})}\left(\frac{r}{R^{\rm th}}\right)^{m}H\!\left(a_{{\rm d}i}\right)$ (61)
$H(a_{{\rm d}i})\!\equiv\!\frac{4\pi-9}{36}\!+\!a_{\rm d}^{2}\frac{\langle x_{\rm d}\rangle}{\Lambda_{\rm d}}\!\left\{\frac{5\pi-8}{30}\hat{a}_{{\rm d}1}^{-4}\!+\!\left(\!\frac{8}{9\pi}+\frac{26+75\pi}{120}\!\right)\!\hat{a}_{{\rm d}1}^{-2}\hat{a}_{{\rm d}2}^{-2}\!-\!\frac{1}{10}\hat{a}_{{\rm d}2}^{-4}\!-\!\frac{1}{5}\hat{a}_{{\rm d}1}^{-2}\hat{a}_{{\rm d}3}^{-3}\!-\!\left(\!\frac{1}{5\pi}+\frac{3\pi}{16}\!\right)\!\hat{a}_{{\rm d}2}^{-2}\hat{a}_{{\rm d}3}^{-2}\!+\!\frac{4}{5}\hat{a}_{{\rm d}3}^{-4}\right\}\!,$ (62)
where we have taken into account the equality $\delta_{\rm d}=(R_{\rm d}/R)^{m}\delta$ with $m=-[(n+3)/2]^{1/2}$ and $R_{\rm d}/R=r/R^{\rm th}$. Taking into account the equality $\hat{a}_{{\rm d}i}^{-2}=\lambda_{{\rm d}i}/\langle x_{\rm d}\rangle$, $H(a_{{\rm d}i})$ adopts the following approximate form as a function of $e_{\rm d}$ (or $\tilde{e}_{\rm d}=-(1-2e_{\rm d})/3$) and $p_{\rm d}$
$H(e_{\rm d},p_{\rm d})\!\approx\!\frac{1}{10}\left(1\!+\!\frac{p_{\rm d}^{-2/3}}{3}\!\left(\!\tilde{e}_{\rm d}^{2/3}\!+\!\tilde{e}_{\rm d}^{8/3}\!-\!\tilde{e}_{\rm d}^{14/3}\!\right)\!\left\{\!1\!+\!\frac{0.6p_{\rm d}\tilde{e}_{\rm d}^{-1}}{1\!+\!\tilde{e}_{\rm d}}\!\left[90\!-\!\left(\frac{1}{\tilde{e}_{\rm d}}\!-\!2\right)^{\!3}\right]\!\right\}\right)~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}$
$\!\approx\!\frac{1}{10}\left[1\!+\!\frac{p_{\rm d}^{-2/3}}{3}\!\left(\!\tilde{e}_{\rm d}^{2/3}\!+\!\tilde{e}_{\rm d}^{8/3}\!-\!\tilde{e}_{\rm d}^{14/3}\right)\!\left(\!1+\!\frac{54p_{\rm d}\tilde{e}_{\rm d}^{-1}}{1+\tilde{e}_{\rm d}}\!\right)\right]\!.~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}$ (63)
Taking into account that $|\tilde{e}_{\rm d}|\la 0.3$, $H(e_{\rm d},p_{\rm d})$ can be further approximated to
$H(e_{\rm d},p_{\rm d})\approx\frac{1}{10}\left\{1+(\tilde{e}_{\rm d}p_{\rm d})^{-1/3}\left[16+\frac{1}{3p_{\rm d}^{4/3}}(\tilde{e}_{\rm d}p_{\rm d})\right]\right\}=\frac{1}{10}\left\{1-3^{1/3}[(1-2e_{\rm d})p_{\rm d}]^{-1/3}\left[16-3^{-2}(1-2e_{\rm d})p_{\rm d}^{-1/3}\right]\right\}.$ (64)
Thus, having performed those two averages, the typical value of e.g. the first Cartesian component of the Lagrangian proptohalo AM takes the form
$J^{\rm L}_{{\rm d}1}\equiv\langle\langle J^{\rm L}_{{\rm d}1}\rangle_{\alpha,\beta,\kappa}\rangle_{\alpha_{\rm d},\beta_{\rm d},\kappa_{\rm d}}\approx\frac{2}{3}\frac{G\bar{\rho}^{1/3}_{0}\delta\,M^{5/3}}{5D(t_{\rm i})}\left(\frac{r}{R^{\rm th}}\right)^{m}H\!\left(e_{\rm d},p_{\rm d}\right)\left(\frac{1-2e}{1+e}\right)^{2/3}p,$ (65)
@@
$J^{\rm L}_{\rm d}=\sqrt{3}J^{\rm L}_{{\rm d}i}\approx\frac{2}{5\times 3^{1/2}D(t_{\rm i})}G\bar{\rho}^{1/3}_{0}\delta\,M^{5/3}\left(\frac{r}{R^{\rm th}}\right)^{m}\,H\!\left(e_{\rm d},p_{\rm d}\right)\left(\frac{1-2e}{1+e}\right)^{2/3}p.$ (66)
Appendix C Ellipticity and prolateness averages
We must average $J^{\rm L}_{\rm d}$ given by equation (66) over the ellipticity and prolateness of the peaks or holes associated with the protohalo and the dominant mass excess or default causing the torque.
Taking into account the following approximate averages of powers of $p_{\rm d}$ from $-e_{\rm d}$ to $e_{\rm d}$ holding for massive haloes, with $e_{\rm d}\la 0.3$ and $p_{\rm d}\ll\sigma_{\rm dp}$,
$\overline{p_{\rm d}^{-1/3}}(e_{\rm d})\approx\frac{6e_{\rm d}^{2/3}}{\sqrt{2\pi}\sigma_{\rm dp}}\,{\rm e}^{-\frac{\bar{p}_{\rm d}^{2}}{2\sigma_{\rm dp}^{2}}}\left[\frac{e_{\rm d}\bar{p}_{\rm d}}{5\sigma_{\rm dp}^{2}}+O(e_{\rm d}^{3})\right]\approx\frac{6e_{\rm d}^{5/3}\bar{p}_{\rm d}}{5\sqrt{2\pi}\sigma_{\rm dp}^{3}}$ (67)
$\overline{p_{\rm d}^{-2/3}}(e_{\rm d})\approx\frac{6e_{\rm d}^{1/3}}{\sqrt{2\pi}\sigma_{\rm dp}}{\rm e}^{-\frac{\bar{p}_{\rm d}^{2}}{2\sigma_{\rm dp}^{2}}}\left[1-\frac{e_{\rm d}^{2}(\bar{p}_{\rm d}^{2}-\sigma_{\rm dp}^{2})}{14\sigma_{\rm dp}^{2}}+O(e_{\rm d}^{3})\right]\approx\frac{6e_{\rm d}^{1/3}}{\sqrt{2\pi}\sigma_{\rm dp}}\left[1+\frac{e_{\rm d}^{2}}{14}\right],$ (68)
@@ Appendix C Ellipticity and prolateness averages
$J^{\rm L}_{\rm d}\approx\frac{2}{5\sqrt{3}}\frac{G\bar{\rho}^{1/3}_{0}\delta\,M^{5/3}}{D(t_{\rm i})}\left(\frac{r}{R^{\rm th}}\right)^{m}\,H(e_{\rm d})\left(\frac{1-2e}{1+e}\right)^{2/3}\bar{p},$ (69)
@@
$H(e_{\rm d})\approx\frac{1}{10}\left[1-\frac{24\times 3^{4/3}}{5\sqrt{2\pi}\sigma_{\rm dp}^{3}}(1-2e_{\rm d})^{-1/3}e_{\rm d}^{5/3}\bar{p}_{\rm d}+\frac{2\times 3^{-2/3}}{\sqrt{2\pi}\sigma_{\rm dp}}(1-2e_{\rm d})^{-2/3}e_{\rm d}^{1/3}\left(1+\frac{e_{\rm d}^{2}}{14}\right)\right],$ (70)
where $\bar{p}$ and $\bar{p}_{\rm d}$ are much smaller than unity, $e$ and $e_{\rm d}$ are smaller than unity and so are also $O(\sigma_{\rm dp})=O(\sigma_{\rm ep})=O(\bar{e}_{\rm d})$. Thus, to leading order in such small quantities we have
$J^{\rm L}_{\rm d}\approx\frac{1}{25\sqrt{3}}\frac{G\bar{\rho}^{1/3}_{0}\delta\,M^{5/3}}{D(t_{\rm i})}\left(1+\frac{2\times 3^{-2/3}}{\sqrt{2\pi}\sigma_{{\rm d}p}}\,e_{\rm d}^{1/3}\right)\left(\frac{r}{R^{\rm th}}\right)^{m}\,\bar{p},$ (71)
where we have kept the term $e_{\rm d}^{1/3}/\sigma_{\rm dp}$ because it is of $O(\bar{e}_{\rm d}^{-2/3})$.
And, taking into account the following approximate average from $-(\bar{e}_{\rm d}+3\sigma_{\rm de})$ to $\bar{e}_{\rm d}+3\sigma_{\rm de}=1.40\bar{e}_{\rm d}$ (even though $e_{\rm d}$ is positive, it is convenient to extend the integral to negative $e_{\rm d}$ values to find the right normalisation constant of the approximately normal PDF)
$\overline{e_{\rm d}^{1/3}}=\frac{6}{\sqrt{2\pi}\sigma_{\rm de}}(1.4\bar{e}_{\rm d})^{4/3}{\rm e}^{-\frac{\bar{e}_{\rm d}^{2}}{2\sigma_{\rm de}^{2}}}\left[\frac{1.4\bar{e}_{\rm d}}{7}\frac{\bar{e}_{\rm d}}{\sigma_{\rm de}^{2}}+O(\bar{e}_{\rm d}^{3})\right]\approx\frac{6\times 1.4^{7/3}}{7\sqrt{2\pi}}\frac{\bar{e}_{\rm d}^{10/3}}{\sigma_{\rm de}^{3}}$ (72)
and the equalities $\sigma_{\rm de}=\bar{e}_{\rm d}/\sqrt{6}$ and $\sigma_{\rm dp}=\bar{e}_{\rm d}/\sqrt{3}$, we find, to leading order, the following average over $e_{\rm d}$
$J^{\rm L}_{\rm d}\approx 0.169\,\frac{G\bar{\rho}^{1/3}_{0}\delta\,M^{5/3}}{D(t_{\rm i})}\left(\frac{r}{R^{\rm th}}\right)^{m}\bar{e}^{-2/3}_{\rm d}\bar{p}=5.07\,\frac{G\bar{\rho}^{1/3}_{0}\delta\,M^{5/3}}{D(t_{\rm i})}\left(\frac{r}{R^{\rm th}}\right)^{m}\bar{e}^{-2/3}_{\rm d}\bar{e}^{4}.$ (73)
Lastly, expressing $\bar{e}$ and $\bar{e}_{\rm d}$ as functions of $\langle x\rangle$ and $\langle x_{\rm d}\rangle$ (see the comment on eq. [12]) and, taking into account that for large masses, as it is the case for the dominant mass excess or default, the average curvature is $\langle x_{\rm d}\rangle\approx\gamma\nu_{\rm d}=(r/R^{\rm th})^{m}\gamma\nu$, where $r/R^{\rm th}$ substantially larger than unity, and adopting $M_{\rm vir}$ masses so that $R_{\rm d}/R\approx r/R^{\rm th}$, we obtain
$J^{\rm L}_{\rm d}\approx 0.347\frac{G\bar{\rho}^{1/3}_{0}\delta\,M^{5/3}}{D(t_{\rm i})}\left(\frac{r}{R^{\rm th}}\right)^{5m/3}\left[\frac{(\gamma\nu)^{1/3}}{\langle x\rangle^{2}+\frac{6}{5}}\right]^{2}.$ (74)